# Aligning AI Research with the Needs of Clinical Coding Workflows: Eight Recommendations Based on US Data Analysis and Critical Review

### v1

@@ 1 Introduction
Clinical coding is a process that transforms clinical notes into a set of alphanumeric codes, which represent diagnoses and procedures during medical visits. This process is essential to tasks like hospital billing and disease prevalence studies. Manual clinical coding is labour-intensive and error-prone Karimi et al. (2017); Li and Yu (2020). To address these issues, automated coding has been widely explored. While many existing studies frame it as a multi-label classification task Mullenbach et al. (2018); Vu et al. (2020); Biswas et al. (2021); Huang et al. (2022); Li et al. (2023), no analysis has examined whether that matches the needs of clinicians.
This position paper critically reviews automated coding studies based on multi-label classification, focusing on those using the public US Medical Information Mart for Intensive Care (MIMIC) datasets Goldberger et al. (2000); Johnson et al. (2016, 2023). For consistency, we use the clinical term ‘code’ instead of ‘label’ throughout this position paper. We focus on the MIMIC datasets as they are far larger than other public datasets. Specifically, MIMIC provides English electronic health records of acute and emergent inpatients, with the latest version, MIMIC-IV, including 331,675 patient admissions. The few other public datasets, such as those released by Pestian et al. (2007) and Miranda-Escalada et al. (2020), contain 978 and 1,000 admissions, respectively. This significant scale difference allows MIMIC to cover a broader range of diagnoses and procedures, making it more suitable for comprehensive analyses.
@@
We then propose new methodologies to support clinical coders beyond automated coding. Given the limited effectiveness of automated coding Edin et al. (2023), manual coding with software assistance remains prevalent, making code auditing essential in clinical workflows to reduce potential human error. Despite their significance, research on AI-based assisted coding and auditing is very limited. This underscores the need for future studies in these areas, as they have the potential to yield practical solutions sooner than relying solely on automated coding.
@@ 2 Clinical Coding Workflow
The implementation of ICD and DRG varies across countries, and the coding inputs can differ depending on healthcare settings (e.g., hospital policies). Previous AI coding studies often use discharge summaries as inputs but have also explored other sources, such as radiology reports Pestian et al. (2007) and multilingual death certificates Névéol et al. (2018). Given all these differences, it is important that AI coding research carefully considers the clinical coding workflow it is addressing. Of note, many references in the following subsection are corporate products, as previous research has largely focused on the automated coding part of the clinical coding workflow.
@@ 2.2 Assisted Coding
In assisted coding, where coders use computer software to manually enter codes, each keystroke, reading, and thought incurs a cost. By minimising the manual effort involved, these tasks become more efficient. This promise of higher efficiency has motivated health tech companies to develop various assisted coding tools.
@@ 2.3 Automated Coding
Automated coding has also been the focus of AI coding research to date, with most studies relying solely on discharge summaries as model inputs. Mullenbach et al. (2018) propose the Convolutional Attention for Multi-Label Classification (CAML) model. CAML combines convolutional neural networks (CNNs) with per-code attention to focus on text sections relevant to each ICD code. This attention mechanism enhances interpretability by highlighting text contributions to the model’s decisions. Li and Yu (2020) extend this framework with the Multi-Filter Residual Convolutional Neural Network (MultiResCNN), which uses multiple filter layers and residual connections for better feature extraction. On the other hand, Vu et al. (2020) introduce the Label Attention Model (LAAT), using a bidirectional long short-term memory (LSTM) to capture clinical context in text. LAAT uses a distinct per-code attention mechanism compared to CAML, incorporating additional steps in attention weight calculations and a hierarchical joint learning strategy, which leads to better predictions for rare ICD codes. Huang et al. (2022) adapt LAAT’s attention mechanism for pretrained language models in ICD coding (PLM-ICD). PLM-ICD also incorporates domain-specific pretraining and segment pooling, addressing challenges such as the domain mismatch between pretraining and clinical text and the large code space.
Recent benchmarking by Edin et al. (2023) compares CAML, MultiResCNN, LAAT, PLM-ICD, and a few other models. PLM-ICD consistently achieves best results on MIMIC-III and MIMIC-IV datasets. However, all models, including PLM-ICD, struggle with rare code, which is a persistent issue in automated coding. Notably, document length had minimal effect on model performance, with little difference when truncating documents from 4,000 to 2,500 words. In Section 3, we demonstrate that prior studies fail to measure key factors critical for real-world applications. Our analysis reveals realistic upper bounds of existing state-of-the-art models compared to human performance.
@@ 3.4 Automated Coding Evaluation Should Match Human Coding Evaluation
Table 2: The code accuracy of PLM-ICD is calculated using the Jaccard Score, with values spread within the 95% confidence interval determined using the Z-score.
@@
The term ‘accuracy’ can be confusing due to its many possible implementations. Different studies have reported human accuracy in clinical coding, but their definitions of accuracy are inconsistent Burns et al. (2012). In this position paper, we will examine two implementations of accuracy. Instance accuracy, or Exact Match Rate (EMR), measures the percentage of instances (medical cases) where the predicted code set exactly matches the true code set. Code accuracy, or Jaccard Score, is defined as the ratio of the size intersection of a predicted code set and a target code set to the size of their union. Instance accuracy is stricter, as it requires a perfect match for every code in an instance (i.e., an episode or medical case).
In automated coding, even if a case contains a single error, human coders must re-code or correct the errors. This means the effort and cost associated with reviewing the entire clinical note are not mitigated, highlighting the need for measuring instance accuracy. On the other hand, code accuracy is useful because not all codes are equally important. When assigning a DRG code, the principal diagnosis code is usually the first and most important determinant factor IHACPA (2023). Given this, when evaluating coding accuracy, it is beneficial to allow for partial matches, acknowledging that capturing overlap in codes can still be valuable.
@@ 3.5 Low Automation Accuracy Suggests Subset-Specific Automation
When considering code accuracy, the median human performance in the UK was 83.2%, with large variance among thirty-two studies (50-98%) Burns et al. (2012). However, the definition of accuracy is inconsistent across the explored studies; some defined inaccurate coding as inaccurate three digit coding, while the majority defined it as inaccurate four digit coding. In fact, many inaccuracies occur at the four digit level Burns et al. (2012) instead of the third digit level. We select PLM-ICD, the best model according to Edin et al. (2023), and report its code accuracy with respect to three-digit, four-digit, and full code levels across three MIMIC splits. For each split, we train a single PLM-ICD instance on the full code prediction task, adjusting the accuracy measure across different digits to account for varying evaluation complexities. Table 2 shows that PLM-ICD’s three digit accuracy, the most generous evaluation measure, is much higher than the other two in all splits. If we compare this result against human accuracy, PLM-ICD is only half as good as an average human at best, indicating there still exists a notable gap to reach full automation.
@@ 4.1.1 The Sequential Task
All three designs add sequential information as part of the input to tackle the code sequencing and dependency issue. Relevant evaluation metrics include Precision@k, the number of steps required to achieve full recall, and the model’s convergence rate with human feedback. Importantly, while human-in-the-loop is not mandatory—since these models can be trained using teacher forcing—the inclusion of human input offers significant benefits. These include higher accuracy over long run, real-time error correction, and improved user trust (important for clinical applications) as users contribute to the model’s learning process.

### v2

@@ 1 Introduction
Clinical coding is a process that transforms clinical notes into a set of alphanumeric codes, which represent diagnoses and procedures during medical visits. This process is essential to tasks like hospital billing and disease prevalence studies. Manual clinical coding is labour-intensive and error-prone Karimi et al. (2017); Li and Yu (2020). To address these issues, automated coding has been widely explored. While many existing studies frame it as a multi-label classification task Mullenbach et al. (2018); Vu et al. (2020); Huang et al. (2022); Li et al. (2023), no analysis has examined whether that matches the needs of clinicians.
This position paper critically reviews automated coding studies based on multi-label classification, focusing on those using the public US Medical Information Mart for Intensive Care (MIMIC) datasets Goldberger et al. (2000); Johnson et al. (2016, 2023). For consistency, we use the clinical term ‘code’ instead of ‘label’ throughout this paper. We focus on the MIMIC datasets as they are far larger than other public datasets. Specifically, MIMIC provides English electronic health records of acute and emergent inpatients, with the latest version, MIMIC-IV, including 331,675 patient admissions. The few other public datasets, such as those released by Pestian et al. (2007) and Miranda-Escalada et al. (2020), contain 978 and 1,000 admissions, respectively. This significant scale difference allows MIMIC to cover a broader range of diagnoses and procedures, making it more suitable for comprehensive analyses.
@@
We then propose new methodologies to support clinical coders rather than automating the entire process. Given the limited effectiveness of automated coding Edin et al. (2023), manual coding with software assistance remains prevalent, with code auditing essential in clinical workflows to reduce errors. Despite its significance, research on AI-assisted coding and auditing is limited. This underscores the need for future studies in these areas, which have the potential to yield practical solutions sooner than fully automated coding.
@@ 2 Clinical Coding Workflow
The implementation of ICD and DRG varies across countries, and the coding inputs can differ depending on healthcare settings (e.g., hospital policies). Previous AI coding studies often use discharge summaries as inputs but have also explored other sources, such as radiology reports Pestian et al. (2007) and multilingual death certificates Névéol et al. (2018). Given all these differences, it is important that AI coding research carefully considers the clinical coding workflow it is addressing. Of note, many references in the following subsection are corporate products, as previous research has largely focused on the automated coding approach in the clinical coding workflow.
@@ 2.2 Assisted Coding
In assisted coding, where coders use computer software to manually enter codes, each keystroke, word read, and thought incurs a cost. By minimising the manual effort involved, these tasks become more efficient. This promise of higher efficiency has motivated health tech companies to develop various assisted coding tools.
@@ 2.3 Automated Coding
Automated coding has also been the focus of AI coding research to date, with most studies relying solely on discharge summaries as model inputs. Mullenbach et al. (2018) propose the Convolutional Attention for Multi-Label Classification (CAML) model. CAML combines convolutional neural networks (CNNs) with per-code attention to focus on text sections relevant to each ICD code. This attention mechanism enhances interpretability by highlighting the text that contributes to the model’s decisions. Li and Yu (2020) extend this framework with the Multi-Filter Residual Convolutional Neural Network (MultiResCNN), which uses multiple filter layers and residual connections for better feature extraction. On the other hand, Vu et al. (2020) introduce the Label Attention Model (LAAT), using a bidirectional long short-term memory (LSTM) to capture clinical context in text. LAAT uses a distinct per-code attention mechanism compared to CAML, incorporating additional steps in attention weight calculations and a hierarchical joint learning strategy, which leads to better predictions for rare ICD codes. Huang et al. (2022) adapt LAAT’s attention mechanism for pretrained language models in ICD coding (PLM-ICD). PLM-ICD also incorporates domain-specific pretraining and segment pooling, addressing challenges such as the domain mismatch between pretraining and clinical text and the large code space.
Recent benchmarking by Edin et al. (2023) compares CAML, MultiResCNN, LAAT, PLM-ICD, and a few other models. PLM-ICD consistently achieves best results on MIMIC-III and MIMIC-IV datasets. However, all models, including PLM-ICD, struggle with rare codes, which is a persistent issue in automated coding. Notably, document length had minimal effect on model performance, with little difference when truncating documents from 4,000 to 2,500 words. In Section 3, we demonstrate that prior studies fail to measure key factors critical for real-world applications. Our analysis reveals realistic upper bounds of existing state-of-the-art models compared to human performance.
@@ 3.4 Automated Coding Evaluation Should Match Human Coding Evaluation
Table 2: The code accuracy of PLM-ICD calculated using Jaccard Score, with values spread within the 95% confidence interval determined using the Z-score.
@@
The term ‘accuracy’ can be confusing due to its many possible implementations. Different studies have reported human accuracy in clinical coding, but their definitions of accuracy are inconsistent Burns et al. (2012). In this position paper, we will examine two implementations of accuracy. Instance accuracy, or Exact Match Rate (EMR), measures the percentage of instances (i.e., episodes of care or medical cases) where the predicted code set exactly matches the true code set. Code accuracy, or Jaccard Score, is defined as the ratio of the size intersection of a predicted code set and a target code set to the size of their union. Instance accuracy is stricter, as it requires a perfect match for every code in an instance.
In automated coding, even if a case contains a single error, human coders must re-code or correct the errors. This means the effort and cost associated with reviewing the entire clinical note are not mitigated, highlighting the need for measuring instance accuracy. On the other hand, code accuracy is useful because not all codes are equally important. When assigning a DRG code, the primary diagnosis code is usually the first and most important determinant factor IHACPA (2023). Given this, when evaluating coding accuracy, it is beneficial to allow for partial matches, acknowledging that capturing overlap in codes can still be valuable.
@@ 3.5 Low Automation Accuracy Suggests Subset-Specific Automation
When considering code accuracy, the median human performance in the UK was 83.2%, with large variance among thirty-two studies (50-98%) Burns et al. (2012). However, the definition of accuracy is inconsistent across the explored studies; some defined inaccurate coding as inaccurate three digit coding, while the majority defined it as inaccurate four digit coding. In fact, many inaccuracies occur at the four digit level Burns et al. (2012) instead of the three digit level. We select PLM-ICD, the best model according to Edin et al. (2023), and report its code accuracy with respect to three-digit, four-digit, and full code levels across three MIMIC splits. For each split, we train a single PLM-ICD instance on the full code prediction task, adjusting the accuracy measure across different digits to account for varying evaluation complexities. Table 2 shows that PLM-ICD’s three digit accuracy, the most generous evaluation measure, is much higher than the other two in all splits. If we compare this result against human accuracy, PLM-ICD is only half as good as an average human at best, indicating there still exists a notable gap to reach full automation.
@@ 4.1.1 The Sequential Task
All three designs add sequential information as part of the input to tackle the code sequencing and dependency issue. Relevant evaluation metrics include Precision@k, the number of steps required to achieve full recall, and the model’s convergence rate with human feedback. Importantly, while it is true that human-in-the-loop is not absolutely mandatory in the suggested designs, the inclusion of human input offers significant benefits. These include higher accuracy over long run, real-time error correction, and improved user trust (critical for clinical applications) as users contribute to the model’s learning process.
@@ 7 Limitations
Acknowledgements
This material is partially supported by the Australian Research Council through a Discovery Early Career Researcher Award and the Commonwealth Scientific and Industrial Research Organisation (CSIRO). We also thank the anonymous reviewers for their constructive feedback on our submission.